4 : 7 = 20 : 35
 4 : 7 = 20 : 35
 4 : 7 = 20 : 35
 4 : 7 = 20 : 35

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The given statement is True.
Consider the two ratio, 4: 7 = 20: 35

= 4/7= 20/35
20/35 is further simplified by dividing both numerator and denominator by 5
we get, = 4/7
Therefore, 4/7 = 4/7


Thanks